# Linear fractional relative risk aversion

### v1

@@ 3.1 Price equilibrium
Some historical remarks.
Lambert (1758) analyzed the equation $Z^{\xi}-Z=\check{x}$, which was tranformed by Euler (1783) into $Z^{\xi}-Z^{\check{\xi}}=(\xi-\check{\xi})xZ^{\xi+\check{\xi}}$. Corless et al. (1996) divide the latter by $\xi-\check{\xi}$ and let $\check{\xi}\to\xi$ to get $\ln Z=xZ^{\xi}$, and then consider the case with $\xi=-1$ and $Z={\rm e}^{W}$ to obtain the Lambert $W$ function defined as $x=W{\rm e}^{W}$. Letting $\tau(Z)$ and $\psi(Z)$ denote some functions of $Z$, we extend Euler’s specification to $(Z/x)^{\xi}\tau(Z)-(Z/x)^{\check{\xi}}=(\xi-\check{\xi})x^{-\check{\xi}}\psi(Z)Z^{\xi+\check{\xi}}$, which implies $(Z/x)^{\xi-\check{\xi}}\tau(Z)-1=(\xi-\check{\xi})\psi(Z)Z^{\xi}$. Setting $\xi=1$, $\check{\xi}=1+\sigma$, $\psi(Z)=1/Z-1$, $\tau(Z)=(1-\beta\sigma)\bigl{[}\frac{1-\beta-Z}{\alpha(1-\beta\sigma)}\bigr{]}^{\beta\sigma}$, and $Z=W$, we obtain $(W/x)^{-\sigma}(1-\beta\sigma)\bigl{[}\frac{1-\beta-W}{\alpha(1-\beta\sigma)}\bigr{]}^{\beta\sigma}-1=-\sigma(1-W)$, which is equivalent to the bottom-right expression of Table 5 with $W=\mu$.
Extending the Lambert $W$ function.
The Lambert $W$ function analyzed in Corless et al. (1996) corresponds to the solution $\mu={\cal W}(x,\alpha,0,0)$ to the first-order condition, $1-\mu=\ln\bigl{(}\frac{\mu}{x}\bigr{)}$ in the top-left expression in Table 5. In the context of monopolistic competition, $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,0,0)}$ corresponds to the equilibrium markup in the CARA model (Behrens et al., 2014, 2017, 2020). We extend the Lambert $W$ function and the markup function in three directions: the direction of $\sigma$, $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,0,\sigma)}$, which is the HARA case; the direction of $\beta$, $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,\beta,0)}$, which is the incomplete gamma case; and the direction of both $\beta$ and $\sigma$, $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,\beta,\sigma)}$, which are the CREMR and hypergeometric cases. [^29] The special case of our generalized Lambert function, ${\cal W}(x,\alpha,\beta,0)$, defined as the solution to the equation in the top-right of Table 5, i.e., $\frac{\mu}{(\frac{1-\beta-\mu}{\alpha})^{\beta}}{\rm e}^{-(1-\beta-\mu)}=x$, is similar to, but differs from $\frac{P(\mu)}{Q(\mu)}{\rm e}^{R(\mu)}=x$ in Mező (2022), where $P(\mu)$, $Q(\mu)$, and $R(\mu)$ are polynomials, because in our analysis the case with $\beta=1$ and $\sigma=0$ is not admissible by Table 4).
HARA case.
As shown in Theorem 2, when $\beta\to 0^{+}$, we get the HARA utility function of the form $u(q)=\frac{K}{\alpha(\sigma-1)}\bigl{[}(1+\alpha\sigma q)^{\frac{\sigma-1}{\sigma}}-1\bigr{]}+C$, where $K>0$ and $C$ are constants. Solving the bottom-left expression of Table 5, we obtain the associated markup function $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,0,\sigma)}$, which nests the quadratic ($\sigma=-1$), log ($\sigma=1$), and CARA ($\sigma\to 0$) utility functions. Imposing $\sigma=-1$, $\sigma=1$, or $\sigma\to 0$ yields the markup function $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,0,-1)}=\frac{1+x}{2x}\geq 1$ for all $0<x\leq 1$, $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,0,1)}=\frac{1}{\sqrt{x}}\geq 1$ for all $0<x\leq 1$, and $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,0,0)}=\frac{1}{W({\rm e}x)}\geq 1$ for all $0<x\leq 1$, respectively, where $W$ is the original Lambert $W$ function. Thus, this case generalizes the Lambert $W$ function in the direction of $\sigma$.
Incomplete gamma case.
As shown in Corollary 1, when $\beta\in(0,1)$ and $\sigma\to 0$, we get the lower incomplete gamma utility of the form $u(q)=\frac{K}{\alpha^{1-\beta}}\gamma(1-\beta,\alpha q)+C$, where $K>0$ and $C$ are constants. Solving the top-right expression of Table 5, we obtain the associated markup function $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,\beta,0)}$, which nests the CARA utility function ($\beta\to 0$). Thus, this case generalizes the Lambert $W$ function in the direction of $\beta$.
CREMR case.
As shown in Theorem 2, when $\beta\to 1^{-}$, we get the CREMR utility of the form $u(q)=\textstyle\frac{K\sigma}{\sigma-1}\frac{(1+\alpha\sigma q)^{1-\frac{1}{\sigma}}}{q}\bigl{[}q-\frac{\sigma-1}{\alpha\sigma}{}_{2}F_{1}\bigl{(}1,1;1+\frac{1}{\sigma};-\frac{1}{\alpha\sigma q}\bigr{)}\bigr{]}+C$, where $K>0$ and $C$ are constants. Recall from Table 4 that $\beta=1$ is possible only in Case 1b, which requires $\alpha>0$ and $\sigma>1$. [^30] Thus, the two special cases of the CREMR given in Table 2, i.e., $u(q)=K\,{\rm Ei}(-\alpha q)$ and $u(q)=K\ln q$, must be excluded from the analysis of monopolistic competition. Solving the bottom-right expression of Table 5 by imposing $\alpha>0$, $\beta=1$, and $\sigma>1$, we obtain the associated markup function
$\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,1,\sigma)}=\frac{1}{1-\frac{1-(1-\sigma)\left[\frac{x}{\alpha(\sigma-1)}\right]^{\sigma}}{\sigma}}=\frac{\sigma}{\sigma-1}\left\{1-\left[\frac{x}{\alpha(\sigma-1)}\right]^{\sigma}\right\}^{-1}>\frac{\sigma}{\sigma-1}>1,$
where $x<\alpha(\sigma-1)$ holds from Table 6.
Hypergeometric case.
The HARA and CREMR cases used in the literature impose either $\beta=0$ or $\beta=1$. The incomplete gamma case allows for intermediate values $\beta\in(0,1)$ but requires that $\sigma\to 0$. We have already shown in Theorem 1 how to deal with more general cases where $\beta\in(0,1)$ and $\sigma\neq 0$, thereby extending the results from the existing literature. As shown in Theorem 1, when $\beta\in(0,1)$ and $\sigma\neq 0$, we get the hypergeometric utility of the form $u(q)=\textstyle\frac{K\sigma}{\sigma-1}(1+\alpha\sigma q)^{\beta-\frac{1}{\sigma}}q^{1-\beta}\left[1+\frac{\beta-\frac{1}{\sigma}}{1-\beta}\,{}_{2}F_{1}\left(1-\frac{1}{\sigma},1;2-\beta;-\alpha\sigma q\right)\right]+C,$ where $K>0$ and $C$ are constants. Solving the bottom-right expression of Table 5, we obtain the associated markup function $\frac{1}{\mu}=\frac{1}{{\cal W}(x,\alpha,\beta,\sigma)}$.
Extending the demand function.
As mentioned above, the first-order condition for utility maximization is given by $(1+\alpha\sigma q)^{\beta-\frac{1}{\sigma}}q^{-\beta}=\frac{\nu p}{K}$, which can be rewritten as $1+\alpha\sigma q=(\frac{\nu p}{K}q^{\beta})^{\frac{1}{\beta-\frac{1}{\sigma}}}=(\frac{\nu}{K}q^{\beta})^{-\frac{\sigma}{1-\beta\sigma}}p^{-\frac{\sigma}{1-\beta\sigma}}$. Hence, the demand function is implicitly given by
$q=\frac{1}{\alpha\sigma}\left(\frac{\nu}{K}q^{\beta}\right)^{-\frac{\sigma}{1-\beta\sigma}}p^{-\frac{\sigma}{1-\beta\sigma}}-\frac{1}{\alpha\sigma},$ (31)
which nests two important cases.
$q=\frac{1}{\alpha\sigma}\left(\frac{\nu}{K}\right)^{-\sigma}p^{-\sigma}-\frac{1}{\alpha\sigma},$ (32)
which is isomorphic to the “translated” CES function in Mrázová and Neary (2017) and the “Pollak family” in Arkolakis et al. (2019). [^31] The demand function in Mrázová and Neary (2017, MN) is given by $q=A^{\rm MN}p^{-\sigma}+B^{\rm MN}$, where $A^{\text{MN}}=\delta$ and $B^{\text{MN}}=\gamma$ (in their notation), whereas that in Arkolakis et al. (2019, ACDR), is given by $q=A^{\rm ACDR}p^{\frac{1}{\gamma}}+B^{\rm ACDR}$, where $A^{\text{ACDR}}=P^{-\frac{1}{\gamma}}$ and $B^{\text{ACDR}}=-\alpha$ (in their notation). It also nests the CARA demand function $q=-\frac{1}{\alpha}\ln\bigl{(}\frac{\nu p}{K}\bigr{)}$ in Behrens and Murata (2007) as seen from letting $\sigma\to 0$.
Second, letting $\beta\to 1^{-}$, we obtain the CREMR demand function
$q=\frac{1}{\alpha\sigma}\left(\frac{\nu}{K}q\right)^{-\frac{\sigma}{1-\sigma}}p^{-\frac{\sigma}{1-\sigma}}-\frac{1}{\alpha\sigma}.$ (33)
Solving this for $p$, we obtain the inverse demand function $p=\frac{K/\nu}{q}(\alpha\sigma q+1)^{\frac{\sigma-1}{\sigma}},$ which is isomorphic to the CREMR inverse demand function in Mrázová et al. (2021). [^32] Our inverse demand function can be rewritten as $p=\frac{K/\nu}{q}(\alpha\sigma)^{\frac{\sigma-1}{\sigma}}(q+\frac{1}{\alpha\sigma})^{\frac{\sigma-1}{\sigma}}$ if $\alpha\sigma$ and $q+\frac{1}{\alpha\sigma}$ are not both negative. This is isomorphic to Mrázová et al.’s (2021) inverse demand function $p=\frac{\beta}{q}(q-\gamma)^{\frac{\sigma-1}{\sigma}}$, where $\beta=(K/\nu)(\alpha\sigma)^{\frac{\sigma-1}{\sigma}}$ and $\gamma=-\frac{1}{\alpha\sigma}$ (in their notation). There are two subcases depending on the sign of $\alpha\sigma$ and $q+\frac{1}{\alpha\sigma}$. First, assume that $\alpha\sigma>0$, which implies $q+\frac{1}{\alpha\sigma}>0$. Then, $(\alpha\sigma)^{\frac{\sigma-1}{\sigma}}$ is real valued, which is consistent with our assumption that $K$ is positive real. Second, assume that $\alpha\sigma<0$, which requires the quantity restriction $q+\frac{1}{\alpha\sigma}>0$ by Corollary 2. Then, $(\alpha\sigma)^{\frac{\sigma-1}{\sigma}}$ is complex, which can be offset by relaxing our assumption of a positive real $K$ and including a complex term in $K$ (see footnote 25).
3.2 Comparative statics results
Proposition 5 (Comparative statics results for LFRRA)
If $\alpha(1-\beta\sigma)>0$ holds (resp., if $\alpha(1-\beta\sigma)<0$ holds), then the markups are decreasing (resp., increasing) in $m$. If either $\alpha=0$ or $\beta=\frac{1}{\sigma}$ holds, then the markups are constant in $m$. The prices are increasing in $m$ and the quantities are decreasing in $m$ regardless of parameter values.
Proof.
@@ 3.2 Comparative statics results
$\frac{{\rm d}\mu}{{\rm d}x}=\frac{\frac{1-\beta\sigma}{x}\left(\frac{\mu}{x}\right)^{-\sigma}\left[\frac{1-\beta-\mu}{\alpha(1-\beta\sigma)}\right]^{\beta\sigma}}{1+\frac{(1-\beta)(1-\mu)}{\alpha\mu}\left(\frac{\mu}{x}\right)^{-\sigma}\left[\frac{1-\beta-\mu}{\alpha(1-\beta\sigma)}\right]^{\beta\sigma-1}}.$ (34)
@@
$\frac{{\rm d}\mu}{{\rm d}x}=\frac{\frac{1-\sigma(1-\mu)}{x}}{1+\frac{(1-\beta)(1-\mu)}{\alpha\mu}\frac{1-\sigma(1-\mu)}{1-\beta\sigma}\frac{\alpha(1-\beta\sigma)}{1-\beta-\mu}}=\frac{\frac{1-\sigma(1-\mu)}{x}}{1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}}.$ (35)
@@
In this case, $\beta=1$ is not admissible. Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold, so that ${1-\beta-\mu}>0$. This, together with $1-\sigma(1-\mu)>0$, implies that the denominator of (35) is positive. Since $x>0$, we know that the numerator of (35) is positive. We thus obtain $\frac{{\rm d}\mu}{{\rm d}x}>0$. Since $x=\frac{\nu m}{K}$, we may conclude that $\frac{{\rm d}\mu}{{\rm d}m}>0$. Consider the subcase with $\beta=0$. In this case, taking the limit of (34) as $\beta\to 0^{+}$ yields $\frac{{\rm d}\mu}{{\rm d}x}=\frac{1}{x}\bigl{[}\frac{1}{\mu}+\bigl{(}\frac{\mu}{x}\bigr{)}^{\sigma}\bigr{]}^{-1}>0$, which implies $\frac{{\rm d}\mu}{{\rm d}m}>0$.
@@
In this case, $\beta=0$ is not admissible, whereas $\sigma>1$ must hold. Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold, so that ${1-\beta-\mu}<0$. This, together with $1-\sigma(1-\mu)<0$, implies that the denominator of (35) is positive. Since $x>0$, we know that the numerator of (35) is negative. We thus obtain $\frac{{\rm d}\mu}{{\rm d}x}<0$. Since $x=\frac{\nu m}{K}$, we may conclude that $\frac{{\rm d}\mu}{{\rm d}m}<0$. Consider the subcase with $\beta=1$. In this case, taking the limit of (34) as $\beta\to 1^{-}$ yields $\frac{{\rm d}\mu}{{\rm d}x}=-\frac{\sigma-1}{x}\bigl{(}\frac{\mu}{x}\bigr{)}^{-\sigma}\bigl{[}\frac{\mu}{\alpha(\sigma-1)}\bigr{]}^{\sigma}<0$, which implies $\frac{{\rm d}\mu}{{\rm d}m}<0$.
@@
In this case, $\beta=0$ is excluded by Step 1 of the proof of Proposition 3 and $\beta=1$ is not admissible by Step 2 of the proof of Proposition 3. Since $\beta\in(0,1)$ in this case, $q>0$ must hold, so that ${1-\beta-\mu}<0$. By the second-order condition in (24) we can show that the denominator of (35) is negative. [^33] Since $\alpha q+\beta>0$ holds for all quantities such that the second-order condition is satisfied (see Step 2 of the proof of Proposition 3 and Table 4), the second-order condition for profit-maximization (24) implies that $\alpha q>-(1-\alpha q-\beta)\frac{\alpha q+\beta}{\alpha\sigma q+1}.$ At any price equilibrium, $1-\mu=\frac{\alpha q+\beta}{\alpha\sigma q+1}$ must hold. This implies $\alpha q>-(1-\alpha q-\beta)(1-\mu)$, so that $\bigl{[}\frac{1-\beta-\mu}{1-\sigma(1-\mu)}\bigr{]}\mu>-(1-\beta)(1-\mu)$, where we use $q=\frac{1}{\alpha}\frac{1-\beta-\mu}{1-\sigma(1-\mu)}$ to get the last inequality. Noting that the left-hand side is negative, we can show that $1<-\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}$, which yields $1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}<0$. Since $x>0$, we know that the numerator of (35) is positive. We thus obtain $\frac{{\rm d}\mu}{{\rm d}x}<0$. Since $x=\frac{\nu m}{K}$, we may conclude that $\frac{{\rm d}\mu}{{\rm d}m}<0$.
@@
In this case, $\beta=0$ and $\beta=1$ are not admissible, whereas $\sigma>1$ must hold. Since $\beta\in(0,1)$ in this case, $q>0$ must hold, so that ${1-\beta-\mu}>0$. By the second-order condition in (24) we can show as before that the denominator of (35) is negative (see footnote 33). Since $x>0$, we know that the numerator of (35) is negative. We thus obtain $\frac{{\rm d}\mu}{{\rm d}x}>0$. Since $x=\frac{\nu m}{K}$, we may conclude that $\frac{{\rm d}\mu}{{\rm d}m}>0$.
@@
where we use $\frac{mx^{\prime}(m)}{x(m)}=1$, which follows from $x=\frac{\nu m}{K}$, to get the last equality. Using (35), we can rewrite it as follows:
$\frac{\partial}{\partial m}\left(\frac{m}{\mu(x(m))}\right)=\textstyle\frac{1}{\mu}\left[1-\frac{x}{\mu}\frac{\frac{1-\sigma(1-\mu)}{x}}{1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}}\right],$ (36)
@@
$\textstyle\frac{\partial}{\partial m}\left(\frac{m}{\mu(x(m))}\right)$ $=$ $\textstyle\frac{1}{\mu}\left[1-\frac{1-\sigma(1-\mu)}{\mu+(1-\beta)(1-\mu)\frac{1-\sigma(1-\mu)}{1-\beta-\mu}}\right]=\textstyle\frac{1}{\mu}\left[\frac{\mu+(1-\beta)(1-\mu)\frac{1-\sigma(1-\mu)}{1-\beta-\mu}-1+\sigma(1-\mu)}{\mu+(1-\beta)(1-\mu)\frac{1-\sigma(1-\mu)}{1-\beta-\mu}}\right]$ (37)
@@
By Table 4, we know that $\frac{1-\mu}{\mu}>0$ except for the special case of Case 1a, where the marginal cost pricing $\frac{1}{\mu}=1$ is admissible when $\beta=0$ (in which case, (37) reduces to $\frac{\partial}{\partial m}\left(\frac{m}{\mu(x(m))}\right)=\frac{1}{\mu+1-\sigma(1-\mu)}$, so that we have $\frac{\partial}{\partial m}\left(\frac{m}{\mu(x(m))}\right)=\frac{1}{2}>0$ at $\mu=1$).
@@
In this case, $\beta=1$ is not admissible. Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold, so that ${1-\beta-\mu}>0$. Since $\beta\in(0,1)$, we know by Table 4 that $\frac{1-\mu}{\mu}>0$. Thus, we focus on the terms in the square brackets of (37). In this case, both the numerator $\frac{1-\beta\sigma}{1-\beta-\mu}$ and the denominator $1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}$ are positive. Hence, $\frac{\partial}{\partial m}\bigl{(}\frac{m}{\mu(x(m))}\bigr{)}>0$. Consider the subcase with $\beta=0$. Taking the limit of (37) as $\beta\to 0^{+}$ yields
@@
In this case, $\beta=0$ is not admissible, whereas $\sigma>1$ must hold. Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold, so that ${1-\beta-\mu}<0$. Since $\beta\in(0,1)$, we know by Table 4 that $\frac{1-\mu}{\mu}>0$. Thus, we focus on the terms in the square brackets of (37). In this case, both the numerator $\frac{1-\beta\sigma}{1-\beta-\mu}$ and the denominator $1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}$ are positive. Hence, $\frac{\partial}{\partial m}\bigl{(}\frac{m}{\mu(x(m))}\bigr{)}>0$. Consider the subcase with $\beta=1$. Taking the limit of (37) as $\beta\to 1^{-}$ yields
@@
In this case, $\beta=0$ is excluded by Step 1 of the proof of Proposition 3 and $\beta=1$ is not admissible by Step 2 of the proof of Proposition 3. Since $\beta\in(0,1)$ in this case, $q>0$ must hold, so that ${1-\beta-\mu}<0$. Since $\beta\in(0,1)$, we know by Table 4 that $\frac{1-\mu}{\mu}>0$. Thus, we focus on the terms in the square brackets of (37). In this case, the numerator $\frac{1-\beta\sigma}{1-\beta-\mu}$ is negative. As shown in footnote 33, the denominator $1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}$ is negative for Case 2a. Hence, $\frac{\partial}{\partial m}\bigl{(}\frac{m}{\mu(x(m))}\bigr{)}>0$.
@@
In this case, $\beta=0$ and $\beta=1$ are not admissible, whereas $\sigma>1$ must hold. Since $\beta\in(0,1)$ in this case, $q>0$ must hold, so that ${1-\beta-\mu}>0$. Since $\beta\in(0,1)$, we know by Table 4 that $\frac{1-\mu}{\mu}>0$. Thus, we focus on the terms in the square brackets of (37). In this case, the numerator $\frac{1-\beta\sigma}{1-\beta-\mu}$ is negative. As shown in footnote 33, the denominator $1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}$ is also negative for Case 2b. Hence, $\frac{\partial}{\partial m}\bigl{(}\frac{m}{\mu(x(m))}\bigr{)}>0$.

### v2

@@ Appendix B.2 Proof of Proposition 4
Appendix B.3 Proof of Proposition 5
@@
$\frac{{\rm d}\mu}{{\rm d}x}=\frac{\frac{1-\beta\sigma}{x}\left(\frac{\mu}{x}\right)^{-\sigma}\left[\frac{1-\beta-\mu}{\alpha(1-\beta\sigma)}\right]^{\beta\sigma}}{1+\frac{(1-\beta)(1-\mu)}{\alpha\mu}\left(\frac{\mu}{x}\right)^{-\sigma}\left[\frac{1-\beta-\mu}{\alpha(1-\beta\sigma)}\right]^{\beta\sigma-1}}.$ (B.6)
@@
$\frac{{\rm d}\mu}{{\rm d}x}=\frac{\frac{1-\sigma(1-\mu)}{x}}{1+\frac{(1-\beta)(1-\mu)}{\alpha\mu}\frac{1-\sigma(1-\mu)}{1-\beta\sigma}\frac{\alpha(1-\beta\sigma)}{1-\beta-\mu}}=\frac{\frac{1-\sigma(1-\mu)}{x}}{1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}}.$ (B.7)
@@
In this case, $\beta=1$ is not admissible. Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold, so that ${1-\beta-\mu}>0$. This, together with $1-\sigma(1-\mu)>0$, implies that the denominator of (B.7) is positive. Since $x>0$, we know that the numerator of (B.7) is positive. We thus obtain $\frac{{\rm d}\mu}{{\rm d}x}>0$. Since $x=\frac{{\lambda}m}{K}$, we may conclude that $\frac{{\rm d}\mu}{{\rm d}m}>0$. Consider the subcase with $\beta=0$. In this case, taking the limit of (B.6) as $\beta\to 0^{+}$ yields $\frac{{\rm d}\mu}{{\rm d}x}=\frac{1}{x}\bigl[\frac{1}{\mu}+\bigl(\frac{\mu}{x}\bigr)^{\sigma}\bigr]^{-1}>0$, which implies $\frac{{\rm d}\mu}{{\rm d}m}>0$.
@@
In this case, $\beta=0$ is not admissible, whereas $\sigma>1$ must hold. Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold, so that ${1-\beta-\mu}<0$. This, together with $1-\sigma(1-\mu)<0$, implies that the denominator of (B.7) is positive. Since $x>0$, we know that the numerator of (B.7) is negative. We thus obtain $\frac{{\rm d}\mu}{{\rm d}x}<0$. Since $x=\frac{{\lambda}m}{K}$, we may conclude that $\frac{{\rm d}\mu}{{\rm d}m}<0$. Consider the subcase with $\beta=1$. In this case, taking the limit of (B.6) as $\beta\to 1^{-}$ yields $\frac{{\rm d}\mu}{{\rm d}x}=-\frac{\sigma-1}{x}\bigl(\frac{\mu}{x}\bigr)^{-\sigma}\bigl[\frac{\mu}{\alpha(\sigma-1)}\bigr]^{\sigma}<0$, which implies $\frac{{\rm d}\mu}{{\rm d}m}<0$.
@@
In this case, $\beta=0$ is excluded by Step 1 of the proof of Proposition 3 and $\beta=1$ is not admissible by Step 2 of the proof of Proposition 3. Since $\beta\in(0,1)$ in this case, $q>0$ must hold, so that ${1-\beta-\mu}<0$. By the second-order condition in (20) we can show that the denominator of (B.7) is negative. [^41] Since $\alpha q+\beta>0$ holds for all quantities such that the second-order condition is satisfied (see Step 2 of the proof of Proposition 3 and Table 3), the second-order condition for profit-maximization (20) implies that $\alpha q>-(1-\alpha q-\beta)\frac{\alpha q+\beta}{\alpha\sigma q+1}.$ At any price equilibrium, $1-\mu=\frac{\alpha q+\beta}{\alpha\sigma q+1}$ must hold. This implies $\alpha q>-(1-\alpha q-\beta)(1-\mu)$, so that $\bigl[\frac{1-\beta-\mu}{1-\sigma(1-\mu)}\bigr]\mu>-(1-\beta)(1-\mu)$, where we use $q=\frac{1}{\alpha}\frac{1-\beta-\mu}{1-\sigma(1-\mu)}$ to get the last inequality. Noting that the left-hand side is negative, we can show that $1<-\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}$, which yields $1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}<0$. Since $x>0$, we know that the numerator of (B.7) is positive. We thus obtain $\frac{{\rm d}\mu}{{\rm d}x}<0$. Since $x=\frac{{\lambda}m}{K}$, we may conclude that $\frac{{\rm d}\mu}{{\rm d}m}<0$.
@@
In this case, $\beta=0$ and $\beta=1$ are not admissible, whereas $\sigma>1$ must hold. Since $\beta\in(0,1)$ in this case, $q>0$ must hold, so that ${1-\beta-\mu}>0$. By the second-order condition in (20) we can show as before that the denominator of (B.7) is negative (see footnote 41). Since $x>0$, we know that the numerator of (B.7) is negative. We thus obtain $\frac{{\rm d}\mu}{{\rm d}x}>0$. Since $x=\frac{{\lambda}m}{K}$, we may conclude that $\frac{{\rm d}\mu}{{\rm d}m}>0$.
@@
where we use $\frac{mx^{\prime}(m)}{x(m)}=1$, which follows from $x=\frac{{\lambda}m}{K}$, to get the last equality. Using (B.7), we can rewrite it as follows:
$\frac{\partial}{\partial m}\left(\frac{m}{\mu(x(m))}\right)=\textstyle\frac{1}{\mu}\left[1-\frac{x}{\mu}\frac{\frac{1-\sigma(1-\mu)}{x}}{1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}}\right],$ (B.8)
@@
$\textstyle\frac{\partial}{\partial m}\left(\frac{m}{\mu(x(m))}\right)$ $=$ $\textstyle\frac{1}{\mu}\left[1-\frac{1-\sigma(1-\mu)}{\mu+(1-\beta)(1-\mu)\frac{1-\sigma(1-\mu)}{1-\beta-\mu}}\right]=\textstyle\frac{1}{\mu}\left[\frac{\mu+(1-\beta)(1-\mu)\frac{1-\sigma(1-\mu)}{1-\beta-\mu}-1+\sigma(1-\mu)}{\mu+(1-\beta)(1-\mu)\frac{1-\sigma(1-\mu)}{1-\beta-\mu}}\right]$ (B.9)
@@
By Table 3, we know that $\frac{1-\mu}{\mu}>0$ except for the special case of Case 1a, where the marginal cost pricing $\frac{1}{\mu}=1$ is admissible when $\beta=0$ (in which case, (B.9) reduces to $\frac{\partial}{\partial m}\left(\frac{m}{\mu(x(m))}\right)=\frac{1}{\mu+1-\sigma(1-\mu)}$, so that we have $\frac{\partial}{\partial m}\left(\frac{m}{\mu(x(m))}\right)=\frac{1}{2}>0$ at $\mu=1$).
@@
In this case, $\beta=1$ is not admissible. Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold, so that ${1-\beta-\mu}>0$. Since $\beta\in(0,1)$, we know by Table 3 that $\frac{1-\mu}{\mu}>0$. Thus, we focus on the terms in the square brackets of (B.9). In this case, both the numerator $\frac{1-\beta\sigma}{1-\beta-\mu}$ and the denominator $1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}$ are positive. Hence, $\frac{\partial}{\partial m}\bigl(\frac{m}{\mu(x(m))}\bigr)>0$. Consider the subcase with $\beta=0$. Taking the limit of (B.9) as $\beta\to 0^{+}$ yields
@@
In this case, $\beta=0$ is not admissible, whereas $\sigma>1$ must hold. Consider the subcase with $\beta\in(0,1)$. In this case, $q>0$ must hold, so that ${1-\beta-\mu}<0$. Since $\beta\in(0,1)$, we know by Table 3 that $\frac{1-\mu}{\mu}>0$. Thus, we focus on the terms in the square brackets of (B.9). In this case, both the numerator $\frac{1-\beta\sigma}{1-\beta-\mu}$ and the denominator $1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}$ are positive. Hence, $\frac{\partial}{\partial m}\bigl(\frac{m}{\mu(x(m))}\bigr)>0$. Consider the subcase with $\beta=1$. Taking the limit of (B.9) as $\beta\to 1^{-}$ yields
@@
In this case, $\beta=0$ is excluded by Step 1 of the proof of Proposition 3 and $\beta=1$ is not admissible by Step 2 of the proof of Proposition 3. Since $\beta\in(0,1)$ in this case, $q>0$ must hold, so that ${1-\beta-\mu}<0$. Since $\beta\in(0,1)$, we know by Table 3 that $\frac{1-\mu}{\mu}>0$. Thus, we focus on the terms in the square brackets of (B.9). In this case, the numerator $\frac{1-\beta\sigma}{1-\beta-\mu}$ is negative. As shown in footnote 41, the denominator $1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}$ is negative for Case 2a. Hence, $\frac{\partial}{\partial m}\bigl(\frac{m}{\mu(x(m))}\bigr)>0$.
@@
In this case, $\beta=0$ and $\beta=1$ are not admissible, whereas $\sigma>1$ must hold. Since $\beta\in(0,1)$ in this case, $q>0$ must hold, so that ${1-\beta-\mu}>0$. Since $\beta\in(0,1)$, we know by Table 3 that $\frac{1-\mu}{\mu}>0$. Thus, we focus on the terms in the square brackets of (B.9). In this case, the numerator $\frac{1-\beta\sigma}{1-\beta-\mu}$ is negative. As shown in footnote 41, the denominator $1+\frac{(1-\beta)(1-\mu)}{\mu}\frac{1-\sigma(1-\mu)}{1-\beta-\mu}$ is also negative for Case 2b. Hence, $\frac{\partial}{\partial m}\bigl(\frac{m}{\mu(x(m))}\bigr)>0$.